A transparent filen of glass of refractive index 1.5 is introduced normally in the path of one of the interfering beams of a Michelson interferometer which is illuminated with a light of wavelength 4800 angstrom. This causes 500 dark fringes to sweep across the field. To determine the thickness of the film.
1
Expert's answer
2019-11-25T10:08:18-0500

We can writer for Michelson Interferometer

without the film

"\u0394_1=2dn_0-2tn_0=m_1 \u03bb (1)"

with the film

"\u0394_2=2ln+2(d-l)n_0-2tn_0=m_2 \u03bb (2)"

where d, t is the Michelson Interferometer arms lengths.

Using (1) and (2) we got

"\u0394= \u0394_1- \u0394_2=2ln-2ln_0=(m_2-m_1)\u03bb (3)"

In our case, n=1.5, n0=1, m2-m1=500, λ=4800 angstrom

l=120×10-6 m


Answer

120×10-6 m